OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation y^2 – 9x^2 + 22y – 72x – 32 = 0?

OpenStudy (helder_edwin):

can u do this one by yourself?

OpenStudy (helder_edwin):

i'll help, if u need it.

OpenStudy (anonymous):

I still kinda need help. I just need to find the center then figure out if its vertical or horizontal right? since its multiple choice? I got stuck but i believe its (-4, -8) and (-4, -14)

OpenStudy (anonymous):

am i right?

OpenStudy (helder_edwin):

i don't know. i have to do the algebra.

OpenStudy (helder_edwin):

\[ \large (y^2+22y+121)-9(x^2+8x+16)=32+121-144 \] \[ \large (y+11)^2-9(x+4)^2=9 \]

OpenStudy (helder_edwin):

so it is vertical and with center (-4,-11)

OpenStudy (anonymous):

okay....

OpenStudy (helder_edwin):

did u understand?

OpenStudy (anonymous):

I think sooo